# font=itshape, raggedright, noindent, leftmargin=0.08in, rightmargin=0.08in, begintext=“ , endtext= ”

### v2

@@ Appendix A Linear Regression with Gaussian Data
We can write $G$ as a quadratic function of $\mathbf{f}$ (fixing $w$) as
$G_{w}(\mathbf{f})$ $=\mathbf{f}^{\top}\left(\tau A^{-1}+\widehat{{\boldsymbol{\xi}}}\widehat{{\boldsymbol{\xi}}}^{\top}\right)\mathbf{f}-2\mathbf{f}^{\top}\left(\tau A^{-1}\overline{\mathbf{x}}-\widehat{{\boldsymbol{\xi}}}\hat{b}+w\hat{b}\right)+\text{const.}$.
$=(\mathbf{f}-\widehat{\mathbf{f}})\left(\tau A^{-1}+\widehat{{\boldsymbol{\xi}}}\widehat{{\boldsymbol{\xi}}}^{\top}\right)(\mathbf{f}-\widehat{\mathbf{f}})+\text{const}.$
Here, $\widehat{\mathbf{f}}$ is calculated as
$\widehat{\mathbf{f}}$ $=\left(\tau A^{-1}+\widehat{{\boldsymbol{\xi}}}\widehat{{\boldsymbol{\xi}}}^{\top}\right)^{-1}\left(\tau A^{-1}\overline{\mathbf{x}}+\widehat{{\boldsymbol{\xi}}}(w-\hat{b})\right)$
$=\left(A_{\tau}-\frac{A_{\tau}\widehat{{\boldsymbol{\xi}}}\widehat{{\boldsymbol{\xi}}}^{\top}A_{\tau}}{1+\widehat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\widehat{{\boldsymbol{\xi}}}}\right)\left(A_{\tau}^{-1}\overline{\mathbf{x}}+\widehat{{\boldsymbol{\xi}}}^{\top}(w-\hat{b})\right),$
@@
$\widehat{\mathbf{f}}=\overline{\mathbf{x}}+A_{\tau}\widehat{{\boldsymbol{\xi}}}(w-\hat{b})-\frac{A_{\tau}\widehat{{\boldsymbol{\xi}}}\widehat{{\boldsymbol{\xi}}}^{\top}\overline{\mathbf{x}}}{1+\widehat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\widehat{{\boldsymbol{\xi}}}}-A_{\tau}\widehat{{\boldsymbol{\xi}}}\frac{\widehat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\widehat{{\boldsymbol{\xi}}}}{1+\widehat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\widehat{{\boldsymbol{\xi}}}}(w-\hat{b}).$
Note that if we denote the predictions of the linear model as $\mathbf{x}_{i}^{\top}\widehat{{\boldsymbol{\xi}}}+\widehat{b}=\hat{y}_{i}$, we can rewrite the above formula as follows:
$\widehat{\mathbf{f}}$ $=\bar{\mathbf{x}}+A_{\tau}\widehat{{\boldsymbol{\xi}}}(w-\widehat{b})\frac{1}{1+\hat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\hat{{\boldsymbol{\xi}}}}-A_{\tau}\hat{{\boldsymbol{\xi}}}\frac{\hat{{\boldsymbol{\xi}}}^{\top}\bar{\mathbf{x}}}{1+\hat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\hat{{\boldsymbol{\xi}}}}$
$=\bar{\mathbf{x}}+A_{\tau}\hat{{\boldsymbol{\xi}}}\frac{(w-\hat{b})}{1+\hat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\hat{{\boldsymbol{\xi}}}}-A_{\tau}\hat{{\boldsymbol{\xi}}}\frac{(\widehat{\overline{\mathbf{y}}}-\hat{b})}{1+\hat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\hat{{\boldsymbol{\xi}}}}$
$=\bar{\mathbf{x}}+A_{\tau}\hat{{\boldsymbol{\xi}}}\frac{w-\widehat{\overline{\mathbf{y}}}}{1+\hat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\hat{{\boldsymbol{\xi}}}}.$
@@
Finally, let’s rewrite $A_{\tau}\widehat{{\boldsymbol{\xi}}}$ and $\widehat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\widehat{{\boldsymbol{\xi}}}$ in terms of interpretable statistical quantities. Recall that $A_{\tau}=\frac{1}{\tau}\left(\frac{X^{\top}X}{N}-\overline{\mathbf{x}}\overline{\mathbf{x}}^{\top}\right)$. Using this, we compute
$A_{\tau}\widehat{{\boldsymbol{\xi}}}$ $=\frac{1}{\tau}\left(\frac{1}{N}\sum_{i=1}^{N}\mathbf{x}_{i}(\underbrace{\mathbf{x}_{i}^{\top}\widehat{{\boldsymbol{\xi}}}}_{=\hat{y}_{i}-\hat{b}})-\overline{\mathbf{x}}\underbrace{\bar{\mathbf{x}}^{\top}\hat{{\boldsymbol{\xi}}}}_{=\frac{1}{N}\sum_{i=1}^{N}\hat{y}_{i}-b}\right)$
$=\frac{1}{N\tau}\sum_{i=1}^{N}(\mathbf{x}_{i}-\overline{\mathbf{x}})(\hat{y}_{i}-\hat{b})$
$=\frac{1}{N\tau}\sum_{i=1}^{N}(\mathbf{x}_{i}-\overline{\mathbf{x}})(\hat{y}_{i}-\hat{\overline{\mathbf{y}}})$
$=\frac{1}{\tau}\operatorname{Cov}(X,\widehat{\mathbf{y}}).$
In the final expression, the term $\operatorname{Cov}(X,\widehat{\mathbf{y}})$ corresponds directly to the previous line. This covariance is a vector that averages data deviations, weighted by prediction deviations. In the line before last, we replaced $\widehat{b}$ with any constant since it is independent of $i$, and the first factor sums to the zero vector. Additionally, we leveraged a key property of linear models: the average of the predictions is the same as the prediction of the average.
@@
$\hat{{\boldsymbol{\xi}}}^{\top}A_{\tau}\hat{{\boldsymbol{\xi}}}=\frac{1}{\tau}\left(\frac{1}{N}\sum_{i=1}^{N}\hat{y}_{i}^{2}-\left(\frac{1}{N}\sum_{i=1}^{N}\hat{y}_{i}\right)^{2}\right)=\frac{1}{\tau}\operatorname{Var}(\widehat{\mathbf{y}}).$
@@
$\Sigma^{-1}=\left(\tau\left(\frac{X^{\top}X}{N}-\bar{\mathbf{x}}\bar{\mathbf{x}}^{\top}\right)^{-1}+\widehat{{\boldsymbol{\xi}}}\widehat{{\boldsymbol{\xi}}}^{\top}\right).$
@@ B.1 Datasets Used in the Experiments
Our experiments are conducted using three numerical datasets and one visual dataset from the literature. The details of the datasets are provided below.
Adult. The Adult dataset (CC BY 4.0 license), derived from the 1994 Census database, comprises 48,842 observations with 14 features, including both continuous and categorical variables [2]. The primary objective is to classify individuals based on whether their annual income exceeds \$50,000 USD. Data preprocessing steps are applied to address missing values and handle categorical features. We applied one-hot encoding to transform the categorical features into a numerical format suitable for our framework.
FICO. The FICO (HELOC) dataset (CC0: Public Domain) consists of home equity line of credit applications submitted by homeowners [7]. It includes 10,459 records with 23 features, comprising both numerical and ordinal variables. The primary objective is to classify applications based on their risk performance, identifying whether an applicant is likely to meet payment obligations or become delinquent. Data preprocessing steps are applied to address missing values.
Housing. The Housing dataset (CC0: Public Domain), sourced from Kaggle, includes information on various house attributes such as lot size, number of rooms, and number of stories [13]. The dataset contains 535 records and 12 features, comprising both numerical and ordinal variables. The primary objective is to predict housing prices based on these features.
MNIST. The MNIST dataset (Data files © Original Authors) is a widely used benchmark in computer vision, consisting of 70,000 grayscale images of handwritten digits (0–9), each represented as a 28×28 pixel matrix [16]. The dataset is divided into 60,000 training samples and 10,000 test samples. The primary objective is to classify images based on the digit they represent. We normalized each of the images to be arrays of shape $(28,28,1)$ with FP32 values in the interval $[0,1]$.
@@ B.2 Experimental Setup
We ran all the experiments on Apple M1 Pro with 32 GB RAM.
For the parameter-sensitive and prediction-risky experiments on the FICO dataset, we trained an MLP with ReLU activation functions and layer widths of $128-32-8-2$. Dropout with a rate of 0.2 was applied after each activation layer to prevent overfitting. The model was trained using a batch size of $128$ for $10,000$ steps.
[…]
@@
Appendix C Langevin Dynamics
We sample data points in this paper by solving a Bayesian optimization problem over distributions on the data space:
$p^{*}=\operatorname*{arg\,min}_{p\in\mathcal{Q}}\mathbb{E}[G]-\tau\mathcal{H}(p).$
where $\mathcal{H}(p)$ denotes the entropy and $\tau>0$ is a temperature parameter. On unconstrained data space, i.e., when $\mathcal{Q}=\mathcal{P}(\mathbb{R}^{d})$, one approach of sampling from $p^{*}$ is to simulate the Langevin dynamics, which is a Stochastic Differential Equation (SDE) given by
$\frac{\mathrm{d}\,X}{\mathrm{d}\,t}=-\frac{1}{\tau}\nabla G(X)+\sqrt{2}\frac{\mathrm{d}\,W}{\mathrm{d}\,t}$
@@ Appendix C Langevin Dynamics
For concrete computation, we work with the standard discretization
$\tilde{\mathbf{x}}_{t+1}=\mathbf{x}_{t}-\eta\nabla G(\mathbf{x}_{t})+\sqrt{2\eta\tau}\varepsilon_{t}$
@@
or otherwise rejected, in which case $\mathbf{x}_{t+1}=\mathbf{x}_{t}$. Here, $\pi\propto e^{-\frac{1}{\tau}G}$ and $q(\mathbf{x}^{\prime}|\mathbf{x})$ is the transition probability of stepping from $\mathbf{x}$ to $\mathbf{x}^{\prime}$ given by $q(\mathbf{x}^{\prime}|\mathbf{x})=\exp(-\frac{1}{4\tau}\|\mathbf{x}^{\prime}-\mathbf{x}+\tau\nabla G(\mathbf{x})\|^{2})$. An implementation of this sampling procedure is provided in the submitted code repository, specifically in the file langevin.py using JAX.
Other Methods and Variants
Beyond the Langevin dynamics approach, several alternative methods exist for sampling data points. For instance, one may use the Picard iteration method or various splitting schemes (see [9]) to discretize and solve the above SDE. Another approach is to restrict the distribution family $\mathcal{Q}$ to be a tractable statistical manifold and apply Variational Inference techniques [8].
It is also worth noting that when sampling over a data manifold, if there exists a mapping $\varphi:\mathbb{R}^{d}\to X$ and $G:X\to\mathbb{R}$, then the corresponding Gibbs distribution on $\mathbb{R}^{d}$ (with respect to the Lebesgue measure) is given by $e^{-\frac{1}{\tau}H}$, where $H=G\circ\varphi:\mathbb{R}^{d}\to\mathbb{R}$. Samples drawn from this distribution can be pushed forward via $\varphi$ to obtain samples in $X$. These pushed-forward samples follow a distribution that can be interpreted as a Gibbs distribution over $X$ with respect to the base measure $\nu$, which is the pushforward of the Lebesgue measure under $\varphi$.
@@ Appendix D Additional Numerical Results
Figure 9 compares the feature distributions between the synthetic dataset generated by our framework and the test data. The box plots represent the range of values for each feature, with blue corresponding to the test data and green representing the generated samples. The Housing dataset [13] used in this experiment contains real estate information such as lot size, number of bathrooms, number of stories, and heating/air conditioning types, aiming to predict house prices. This figure provides a clear visualization of how the generated data differs from the test data in terms of feature distributions. For instance, as the number of bathrooms and stories increases, the model predictions diverge. Additionally, hot water heating and air conditioning exhibit a distinct concentration in the synthetic data, with most generated samples clustering around higher values compared to the test data. This suggests that these features play a prominent role in distinguishing instances where the models behave differently. Overall, this figure offers insights into how the generated samples differ from the original dataset, highlighting key feature distributions that drive divergence in model predictions and providing a deeper understanding of how our framework probes model behavior.
Prediction-risky samples. To further investigate data samples near the decision boundary, we present the distributions of all features in the original dataset and the generated prediction-risky samples in Figure 10. These density plots provide a comprehensive view of the differences between the generated samples and the original data across multiple features. By analyzing these distributions, we can observe how the model identifies borderline cases based on different financial attributes. Across multiple features, the generated prediction-risky samples exhibit a much narrower distribution compared to the original data. This suggests that the model focuses on a specific subset of feature values when identifying borderline cases.

### v3

@@ Appendix A Linear Regression with Gaussian Data
[…]
Note that if we denote the predictions of the linear model as $\mathbf{x}_{i}^{\top}\widehat{{\bm{\xi}}}+\widehat{b}=\hat{y}_{i}$, we can rewrite the above formula as follows:
$\widehat{\mathbf{f}}$ $=\bar{\mathbf{x}}+A_{\tau}\widehat{{\bm{\xi}}}(w-\widehat{b})\frac{1}{1+\hat{{\bm{\xi}}}^{\top}A_{\tau}\hat{{\bm{\xi}}}}-A_{\tau}\hat{{\bm{\xi}}}\frac{\hat{{\bm{\xi}}}^{\top}\bar{\mathbf{x}}}{1+\hat{{\bm{\xi}}}^{\top}A_{\tau}\hat{{\bm{\xi}}}}$
$=\bar{\mathbf{x}}+A_{\tau}\hat{{\bm{\xi}}}\frac{(w-\hat{b})}{1+\hat{{\bm{\xi}}}^{\top}A_{\tau}\hat{{\bm{\xi}}}}-A_{\tau}\hat{{\bm{\xi}}}\frac{(\widehat{\overline{\mathbf{y}}}-\hat{b})}{1+\hat{{\bm{\xi}}}^{\top}A_{\tau}\hat{{\bm{\xi}}}}$
$=\bar{\mathbf{x}}+A_{\tau}\hat{{\bm{\xi}}}\frac{w-\widehat{\overline{\mathbf{y}}}}{1+\hat{{\bm{\xi}}}^{\top}A_{\tau}\hat{{\bm{\xi}}}}$.
@@
Finally, let’s rewrite $A_{\tau}\widehat{{\bm{\xi}}}$ and $\widehat{{\bm{\xi}}}^{\top}A_{\tau}\widehat{{\bm{\xi}}}$ in terms of interpretable statistical quantities. Recall that $A_{\tau}=\frac{1}{\tau}\left(\frac{X^{\top}X}{N}-\overline{\mathbf{x}}\overline{\mathbf{x}}^{\top}\right)$. Using this, we compute
$A_{\tau}\widehat{{\bm{\xi}}}$ $=\frac{1}{\tau}\left(\frac{1}{N}\sum_{i=1}^{N}\mathbf{x}_{i}(\underbrace{\mathbf{x}_{i}^{\top}\widehat{{\bm{\xi}}}}_{=\hat{y}_{i}-\hat{b}})-\overline{\mathbf{x}}\underbrace{\bar{\mathbf{x}}^{\top}\hat{{\bm{\xi}}}}_{=\frac{1}{N}\sum_{i=1}^{N}\hat{y}_{i}-b}\right)$
[…]
@@
[…]
@@
$\Sigma^{-1}=\left(\tau\left(\frac{X^{\top}X}{N}-\bar{\mathbf{x}}\bar{\mathbf{x}}^{\top}\right)^{-1}+\widehat{{\bm{\xi}}}\widehat{{\bm{\xi}}}^{\top}\right).$
@@ B.1 Datasets Used in the Experiments
Our experiments are conducted using four numerical datasets and one visual dataset from the literature. The details of the datasets are provided below.
Adult. The Adult dataset (CC BY 4.0 license), derived from the 1994 Census database, comprises 48,842 observations with 14 features, including both continuous and categorical variables (adult_data). The primary objective is to classify individuals based on whether their annual income exceeds \$50,000 USD. Data preprocessing steps are applied to address missing values and handle categorical features. We applied one-hot encoding to transform the categorical features into a numerical format suitable for our framework.
FICO. The FICO (HELOC) dataset (CC0: Public Domain) consists of home equity line of credit applications submitted by homeowners (fico_data). It includes 10,459 records with 23 features, comprising both numerical and ordinal variables. The primary objective is to classify applications based on their risk performance, identifying whether an applicant is likely to meet payment obligations or become delinquent. Data preprocessing steps are applied to address missing values.
Housing. The Housing dataset (CC0: Public Domain), sourced from Kaggle, includes information on various house attributes such as lot size, number of rooms, and number of stories (housing_data). The dataset contains 535 records and 12 features, comprising both numerical and ordinal variables. The primary objective is to predict housing prices based on these features.
MNIST. The MNIST dataset (Data files © Original Authors) is a widely used benchmark in computer vision, consisting of 70,000 grayscale images of handwritten digits (0–9), each represented as a 28×28 pixel matrix (mnist_data). The dataset is divided into 60,000 training samples and 10,000 test samples. The primary objective is to classify images based on the digit they represent. We normalized each of the images to be arrays of shape $(28,28,1)$ with FP32 values in the interval $[0,1]$.
Wine Recognition Dataset. The Wine Recognition dataset contains the results of chemical and physical analyses of wines produced in three regions of Italy (wine). Each record is described by 13 continuous features obtained through chemical analysis, including alcohol content, malic acid, magnesium, color intensity, and hue. The primary objective is to classify wines according to their region of origin.
@@ B.2 Experimental Setup
We ran all the experiments on a personal Apple M1 Pro with 32 GB RAM. In our experiments compute time was not a significant factor, all experiments were completed within minutes.
For the parameter-sensitive and prediction-risky experiments on the FICO dataset, we trained an MLP with ReLU activation functions and layer widths of 1283282128-32-8-2. Dropout with a rate of 0.2 was applied after each activation layer to prevent overfitting. The model was trained using a batch size of $128$ for $10,000$ steps.
For the image experiments, we used an MLP with layer widths of 10241024-128128-$10$, where each layer included a ReLU activation, followed by a dropout layer with a rate of 0.2. The CNN architecture consisted of two convolutional blocks with feature sizes 326432-64. Each block followed the structure:  $\text{ Conv}\rightarrow\text{ReLU}\rightarrow\text{Conv}\rightarrow\text{ReLU}\rightarrow\text{ max\_pool}\rightarrow\text{Dropout}$, where the convolutional kernels had a size of $3\times 3$, the max pooling window was $2\times 2$, and the dropout rate was 0.2.
@@
Appendix C Langevin Dynamics solving the Bayesian Learning Problem
In solving equation 3, we sample data points by solving a Bayesian optimization problem over distributions on the data space:
$p^{*}=\operatorname*{arg\,min}_{p\in\mathcal{Q}}\mathbb{E}[G]-\tau\mathcal{H}(p),$
where $\mathcal{H}(p)$ denotes the entropy and $\tau>0$ is a temperature parameter.
The intuition behind this objective is that instead of minimizing a function $G$ we find a distribution over the space $X$ which balances having its distributional mass focused on points where $G$ is low and also a spread widely. This objective is also called the Evidence Lower BOund (ELBO) and interpreting the loss as negative log-likelihood the solution can be interpreted as Bayes’s formula updating beliefs upon observing new data, as shown by Zellner zellner1988optimal.
C.1 Metropolis Hastings Langevin Algorithm (MALA)
On unconstrained data space, i.e., when $\mathcal{Q}=\mathcal{P}(\mathbb{R}^{d})$, one approach of sampling from $p^{*}$ is to simulate the Langevin dynamics, which is a Stochastic Differential Equation (SDE) given by
$\frac{\mathrm{d}\,X}{\mathrm{d}\,t}=-\frac{1}{\tau}\nabla G(X)+\sqrt{2}\frac{\mathrm{d}\,W}{\mathrm{d}\,t},$
@@ C.1 Metropolis Hastings Langevin Algorithm (MALA)
For concrete computation, we work with the standard Euler-Maruyama discretization
$\tilde{\mathbf{x}}_{t+1}=\mathbf{x}_{t}-\eta\nabla G(\mathbf{x}_{t})+\sqrt{2\eta\tau}\varepsilon_{t},$
@@
or otherwise rejected, in which case $\mathbf{x}_{t+1}=\mathbf{x}_{t}$. Here, $\pi\propto e^{-\frac{1}{\tau}G}$ and $q(\mathbf{x}^{\prime}|\mathbf{x})$ is the transition probability of stepping from $\mathbf{x}$ to $\mathbf{x}^{\prime}$ given by $q(\mathbf{x}^{\prime}|\mathbf{x})=\exp(-\frac{1}{4\tau}\|\mathbf{x}^{\prime}-\mathbf{x}+\tau\nabla G(\mathbf{x})\|^{2})$. An implementation of this sampling procedure is provided in the submitted code repository, specifically in the file langevin.py using JAX/PyTorch and Numpy.
The temperature hyperparameter $\tau$ determines the balance between exploration of the data space versus minimizing $G$. In the figure below we show the effect of changing this parameter (as well as the effect of changing the tunable parameter of the regularizing term).
Figure 8: The probing functions in Figure 2, where in the first there columns the data generation in the second line trained with different temperatures. The last column shows the effect of a $\ell_{2}$ regularizing term (with anchor point shown in orange) with a smaller strength.
C.2 Smoothing
The acceptance/rejection is the critical step to ensure convergence of the trajectory to the limiting Gibbs-Boltzmann distribution $\propto e^{-\frac{1}{\tau}G}$. The drift term $\nabla G(\mathbf{x})$ speeds up the convergence by ensuring that more of the proposals $\widetilde{\mathbf{x}}$ will be accepted since this becomes (noisy) gradient descent.
In case the function $G$ is locally flat, such as those functions created from tree-based models (Decision Tree, Random Forest, XGBoost) then the gradient term is always $\mathbf{0}$, reducing the proposals to simple random walk. This is called the Metropolis Hastings (MH) algorithm and it also has the same limiting distribution, and the transition probabilities in the acceptance ratio cancel since $q(\mathbf{x}|\mathbf{x}^{\prime})=q(\mathbf{x}^{\prime}|\mathbf{x})$.
However in this case at any step the proposals are random and the point $\mathbf{x}$ does not see if it is near a decision boundary or not. For this purpose we propose using the gradients of the smoothed function $G_{s}(\mathbf{m})=\int G(\mathbf{x})\mathcal{N}(\mathbf{m},\sigma)(\mathbf{x})\mathrm{d}\mathbf{x}$ for the proposals. With a larger $\sigma$, the point $\mathbf{x}$ “sees” a wider horizon for its proposals. Indeed the gradient of $G_{s}$ is
$\nabla{G}_{s}(\mathbf{x})=\int G(\mathbf{x}+\sigma\mathbf{\varepsilon})\mathbf{\varepsilon}\mathcal{N}(\mathbf{0},I)(\mathbf{\varepsilon})\mathrm{d}\mathbf{\varepsilon}\approx\frac{1}{J}\sum_{j=1}^{J}G(\mathbf{x}+\sigma\mathbf{\varepsilon}_{j})\mathbf{\varepsilon}_{j},\qquad\text{with }\mathbf{\varepsilon}_{j}\sim\mathcal{N}(\mathbf{0},I).$
In our implementation we applied proposals with $\nabla G_{s}$ and acceptance/rejection with $G$ In the code, we also assumed $\nabla G_{s}(\mathbf{x})=\nabla G_{s}(\mathbf{x}^{\prime})$ for computational simplicity in calculating the transition probabilities $q(\mathbf{x}^{\prime}|\mathbf{x})$. This is approximately true when $G$ is a step function and the step size $\eta$ is small.
C.3 Other Methods and Variants
Beyond the Langevin dynamics approach, several alternative methods exist for sampling data points. For instance, one may use the Picard iteration method or various splitting schemes (see (picarditer)) to discretize and solve the above SDE. Another approach is to restrict the distribution family $\mathcal{Q}$ to be a tractable statistical manifold and apply Variational Inference techniques (gangulyVI).
It is also worth noting that when sampling over a data manifold, if there exists a mapping $\varphi:\mathbb{R}^{d}\to\mathcal{X}$ and $G:\mathcal{X}\to\mathbb{R}$, then the corresponding Gibbs distribution on $\mathbb{R}^{d}$ (with respect to the Lebesgue measure) is given by $e^{-\frac{1}{\tau}H}$, where $H=G\circ\varphi:\mathbb{R}^{d}\to\mathbb{R}$. Samples drawn from this distribution can be pushed forward via $\varphi$ to obtain samples in $\mathcal{X}$. These pushed-forward samples follow a distribution that can be interpreted as a Gibbs distribution over $\mathcal{X}$ with respect to the base measure $\nu$, which is the pushforward of the Lebesgue measure under $\varphi$.
@@ Appendix D Additional Numerical Results
Figure 10 compares the feature distributions between the synthetic dataset generated by our framework and the test data. The box plots represent the range of values for each feature, with blue corresponding to the test data and green representing the generated samples. The Housing dataset (housing_data) used in this experiment contains real estate information such as lot size, number of bathrooms, number of stories, and heating/air conditioning types, aiming to predict house prices. This figure provides a clear visualization of how the generated data differs from the test data in terms of feature distributions. For instance, as the number of bathrooms and stories increases, the model predictions diverge. Additionally, hot water heating and air conditioning exhibit a distinct concentration in the synthetic data, with most generated samples clustering around higher values compared to the test data. This suggests that these features play a prominent role in distinguishing instances where the models behave differently. Overall, this figure offers insights into how the generated samples differ from the original dataset, highlighting key feature distributions that drive divergence in model predictions and providing a deeper understanding of how our framework probes model behavior.
Prediction-risky samples.
This experiment explores data samples near the model’s decision boundary, where predictions are inherently uncertain. To guide this analysis, we pose the question:
The density plots in Figure 11 compare the distributions of two representative features in the original data and the generated samples. The feature NumTrades60Ever2DerogPubRec represents the number of past credit trades with payments delayed by at least 60 days, serving as a key indicator of delinquency. As shown in the figure, the distribution of risky samples follows the original data closely in the lower range but exhibits a stronger peak around zero. This suggests that the model considers individuals with few or no past delinquencies as borderline cases, likely due to the absence of strong negative or positive indicators, making classification more uncertain. The feature MSinceOldestTradeOpen indicates the number of months since a customer’s first credit line was opened, capturing the length of their credit history. As shown in Figure 11, the distribution of risky samples is sharply concentrated around 400 months ($\sim$33 years), whereas the original data spans a broader range. This suggests that the model associates long credit histories with greater uncertainty. The pronounced peak reflects how the model treats long-established credit profiles as ambiguous when making predictions.
To further investigate data samples near the decision boundary, we present the distributions of all features in the original dataset and the generated prediction-risky samples in Figure 12. These density plots provide a comprehensive view of the differences between the generated samples and the original data across multiple features. By analyzing these distributions, we can observe how the model identifies borderline cases based on different financial attributes. Across multiple features, the generated prediction-risky samples exhibit a much narrower distribution compared to the original data. This suggests that the model focuses on a specific subset of feature values when identifying borderline cases.
@@
Table 1: Characteristics of the wine recognition dataset and the generated data
Characteristic Class 0 Class 1 Class 2 Generated
Samples 59 71 48 50
Alcohol $13.74_{\pm 0.63}$ $12.28_{\pm 0.53}$ $13.15_{\pm 0.52}$ $12.75_{\pm 0.61}$
Malic acid $2.01_{\pm 0.68}$ $1.93_{\pm 1.00}$ $3.33_{\pm 1.08}$ $1.59_{\pm 0.46}$
Ash $2.46_{\pm 0.22}$ $2.45_{\pm 0.31}$ $2.44_{\pm 0.18}$ $1.83_{\pm 0.25}$
Alcalinity of ash $17.04_{\pm 2.52}$ $20.24_{\pm 3.33}$ $21.42_{\pm 2.23}$ $19.37_{\pm 4.95}$
Magnesium $106.34_{\pm 10.41}$ $4.55_{\pm 16.63} 99$$9.31_{\pm 10.78}$ $98.9_{\pm 28.91}$
Total phenols $84_{\pm 0.34} 2.2$$26_{\pm 0.54}$ $1.68_{\pm 0.35}$ $1.69_{\pm 0.43}$
Flavanoids $2.98_{\pm 0.39}$ $2.08_{\pm 0.7}$ $0.78_{\pm 0.29}$ $2.11_{\pm 0.82}$
Nonflavanoid phenols $0.29_{\pm 0.07}$ $0.36_{\pm 0.12}$ $0.45_{\pm 0.12}$ $0.3_{\pm 0.08}$
Proanthocyanins $1.90_{\pm 0.4}$ $63_{\pm 0.6} 1.1$$15_{\pm 0.4}$ $2.46_{\pm 0.41}$
Color intensity $5.53_{\pm 1.23}$ $3.09_{\pm 0.92}$ $7.4_{\pm 2.3}$ $2.34_{\pm 0.6}$
Hue $1.06_{\pm 0.16}$ $1.06_{\pm 0.2}$ $0.68_{\pm 0.11}$ $1.27_{\pm 0.14}$
OD280/OD315 $3.16_{\pm 0.35}$ $2.78_{\pm 0.49}$ $1.68_{\pm 0.27}$ $1.69_{\pm 0.39}$
Proline $1115.71_{\pm 221.64}$ $519.5_{\pm 156.1}$ $629.9_{\pm 113.9}$ $1200.8_{\pm 233.3}$
@@
Additionally, we compare the counterfactual distribution produced by our probing scenario with counterfactual examples generated by DiCE for a representative Adult dataset instance. In this experiment, the factual instance represents a North-American White Female, predicted to earn less than \$50K. Figure 15 shows the marginal feature distributions for DiCE counterfactuals, our generated counterfactual samples, and the original training data, together with the factual instance. As expected, DiCE produces a finite set of discrete counterfactuals with relatively high variance across many features, and often explores extreme or low-density regions of the input space. In contrast, our method generates a full stationary distribution over counterfactual inputs, producing tightly clustered samples that remain stable across independent runs. The resulting distribution concentrates on plausible, high-density regions of the dataset and alters only the features necessary for flipping the model’s decision. We note that all generated samples strictly satisfy the decision-flipping condition. This highlights the complementary nature of the two approaches. DiCE provides diverse point-wise counterfactuals, while our probing scenario characterizes the underlying counterfactual landscape by modeling the distribution of model-aligned inputs.
Figure 15: Comparison between our generated counterfactual distribution and DiCE counterfactual examples for the same factual instance.